# Edge densities of drawings of graphs with one forbidden cell

### v2

@@ 6.3 Non-Quasiplanar [Uncaptioned image]-Free Simple Drawings
There are infinitely many graphs that admit simple [Uncaptioned image]-free drawings but no non-homotopic quasiplanar drawings.
Proof 6.6.
Start with an arbitrary graph $G$ that does not admit a non-homotopic quasiplanar drawing (any graph on $n$ vertices and >8n20>8n-20 edges suffices). Let $\Gamma$ be any simple drawing of $G$. If $\Gamma$ contains no [Uncaptioned image]-cells, we are done. Otherwise, let $c$ be a [Uncaptioned image]-cell and $e=uv$ one of the edges on its boundary. Add a new vertex $w$ in $c$ to $\Gamma$ and draw an edge $uw$ that starts at $u$, follows $e$ closely (such that $uw$ crosses the exact same edges as $e$ does between $u$ and the part where $e$ bounds the [Uncaptioned image]-cell containing $w$, in the same order) until entering $c$ and ending in $w$; see Figure˜8. In the new drawing $\Gamma^{+}$, the corridor bounded by $e$ and $uw$ only contains a [Uncaptioned image]-cell and some [Uncaptioned image]-cells. Moreover, $c$ becomes a (degenerate) [Uncaptioned image]-cell in $\Gamma^{+}$. All other cells in $\Gamma^{+}$ incident to $uw$ have a corresponding cell of the same type in $\Gamma$ with the only difference that the edge segments of $e$ are replaced by edge segments of $uw$. Therefore, adding $w$ and $uw$ reduces the number of [Uncaptioned image]-cells. Further, the new drawing $\Gamma^{+}$ retains simplicity, as $uw$ crosses a subset of the edges crossed by $e$.
Repeating this process, a [Uncaptioned image]-free simple drawing $\Gamma^{\prime}$ of a graph $G^{\prime}$ with $G\subseteq G^{\prime}$ is obtained. Since $G$ does not admit a non-homotopic quasiplanar drawing, neither does $G^{\prime}$, and so, $G^{\prime}$ is as desired.
Figure 8: Illustration of Proposition˜6.5. By adding a vertex and an edge, the number of Refer to caption-cells in a drawing can be reduced by one.
@@ 7.1 Simple [Uncaptioned image]-Free Drawings
We begin with the construction of [Uncaptioned image]-free simple drawings on the torus with a quadratic number of edges.
@@
Now, we add for each vertex $v$ in column $j$ and each of the $k$ slopes $s\in S_{j}$ an edge of that slope $s$ from $v$ to a vertex in column $j+3$. That is, for each vertex $v_{i,j}$, and for each slope $s\in S_{j}$, we add an edge with slope $s$ between $v_{i,j}$ and $v_{i+3s,j+3}$ that passes through $v_{i+s,j+1}$ and $v_{i+2s,j+2}$. See Figure˜9. Note that this is not yet a valid drawing by our definition, as edges pass through vertices and there are points in which more than two edges cross.
To fix this, consider any vertex $v_{i,j}$. There are $k$ edges incident to $v_{i,j}$ that end in column $j+3$, and $k$ edges incident to $v_{i,j}$ that start in column j3j-3. Since $S_{j}=S_{j-3}=S_{j+3}$, both these sets of $k$ edges have the same $k$ slopes. Additionally, there is a set $A$ of $2k$ edges that pass through $v_{i,j}$ but are not incident to $v_{i,j}$. The $2k$ edges in $A$ have the $2k$ slopes in $\{0,\ldots,3k-1\}-S_{j}$. The situation in a small disk around $v_{i,j}$ is depicted in Figure˜10 (Left). We perturb the $2k$ edges of $A$ slightly within the disk around $v_{i,j}$ such that locally we obtain a valid drawing with the additional property that each of the $2k$ edges emanating from $v_{i,j}$ has its first crossing with a different of the $2k$ edges in $A$ (see Figure˜10 (Right)). This ensures that none of the cells incident to $v_{i,j}$ is a [Uncaptioned image]-cell. Repeating this procedure at every vertex $v_{i,j}$, the resulting drawing is [Uncaptioned image]-free. If after these local changes around vertices, other points in which more than two edges cross remain, locally perturb these as well to obtain a valid drawing $\Gamma(k)$.
Figure 10: Illustration for Construction˜9. In a drawing, when a set $S$ of $2k$ edges intersect at a vertex of degree $2k$, the edges can be locally redrawn without introducing new edge intersections such that the vertex is not incident to any Refer to caption-cells.
Finally, we argue that $\Gamma(k)$ is a simple drawing. Note that any edge in $\Gamma(k)$ spans (including its endpoints) at most $9k$ rows and exactly $4$ columns. Hence the size of the grid ensures that two edges do not intersect multiple times by wrapping around the torus. Further, before the perturbations, all edges behave like straight lines locally, and hence any two (unperturbed) edges intersect at most once. Because the perturbations do not introduce any new crossings, $\Gamma(k)$ is simple.
Based on the toroidal construction, we obtain simple [Uncaptioned image]-free drawings in the plane with $\Omega(n^{3/2})$ edges. While this leaves a significant gap to the trivial upper bound of $\binom{n}{2}$, we believe that there are [Uncaptioned image]-free drawings with $\Omega(n^{2})$ edges, see ˜9.2.
@@
Proof 7.2.
Let $k$ be some positive integer. Start with the toroidal drawing $\Gamma(k)$ described in Construction˜9 but with an underlying $20k\times 9k$-grid instead of a $20k\times 9$-grid. See again Figure˜9 for a visualization. Now, since we cannot place $\Gamma(k)$ in the plane as is, we cut the torus between two neighboring columns (say between column 9k19k-1 and 0). The cut is positioned so that it does not pass through any edge crossings. The resulting “drawing” can be embedded in the plane but contains many halved edges that end on the boundaries of the cut we performed. We take care of these half-edges by placing a new vertex each at the point in the plane where the half-edge ends, and, in doing so, add degree 1-vertices.
@@ 7.2 Non-Homotopic [Uncaptioned image]-Free Drawings
For each integer $k\geq 2$, we construct a simple $k$-regular graph $G(k)$ on $n=3\cdot(3k-1)$ vertices together with a simple [Uncaptioned image]-free drawing; see Figures˜11 and 12 for an illustration. The graph $G(k)$ consists of three isomorphic copies of a graph $G^{\prime}(k)$ on 3k13k-1 vertices. To describe $G^{\prime}(k)$ (and its drawing), we start with a set $A$ of 3k13k-1 equidistant points on a circle $C_{\rm out}$. The points in $A$ will be the vertices of $G^{\prime}(k)$, which we denote $a_{0},\ldots,a_{3k-2}$ in cyclic ordering around $C_{\rm out}$. For each $i=0,\ldots,3k-2$ draw a straight-line edge $e_{i}$ between $a_{i}$ and $a_{i+k}$ (throughout this construction all indices are taken modulo 3k13k-1). Let $b_{i}$ be the crossing point of $e_{i}$ and $e_{i+1}$, and $B$ the set of all such $b_{i}$.
@@
Figure 11: Illustration of the $k$-regular graph $G^{\prime}(k)$ on 3k13k-1 vertices for $k=2$ and $k=5$. At each of the 3k13k-1 points in $B$ (blue disks), $k$ edges are pairwise crossing. Corridors are indicated in red. (Due to space reasons, only some corridors are shown for $k=5$.)
@@
Then, $b_{0},\ldots,b_{3k-2}$ are equidistant points (blue disks in Figure˜11) on a smaller circle $C_{\rm in}$ concentric to $C_{\rm out}$. Note that $e_{i}$ crosses $e_{i+1}$ at $b_{i}$ and $e_{i-1}$ at $b_{i-1}$. Next, we add further edges such that each $a_{i}$ has exactly $k$ incident edges and at each $b_{i}$, exactly $k$ edges pairwise cross. To this end, we insert edges from $a_{i}$ to $a_{i+k+1},\ldots,a_{i+\lfloor 3k/2\rfloor}$ where the edge $a_{i}a_{i+k+j}$ with $j=1,\ldots,\lfloor k/2\rfloor$ goes through $b_{i-j-1}$ and $b_{i+2j}$; see again Figure˜11 for an illustration. Let us emphasize (as we will need it later) that every edge passes through exactly two points in $B$. This completes the construction of $G^{\prime}(k)$. Observe that the described drawing of $G^{\prime}(k)$ is simple (in fact, almost straight-line) and every vertex of $G^{\prime}(k)$ lies on the same cell.
@@
Figure 12: Connecting the drawings of $G^{\prime}_{1}(k)$, $G^{\prime}_{2}(k)$, and $G^{\prime}_{3}(k)$ to a Refer to caption-free drawing of $G(k)$. The center of each $G^{\prime}_{i}(k)$ is depicted as a yellow disk. The two edges $f$ and $f^{\prime}$ cross twice but the lens they create is non-empty, since it contains the vertices $u$ and $v$.
@@
We now take three copies of $G^{\prime}(k)$, called $G^{\prime}_{1}(k),G^{\prime}_{2}(k),G^{\prime}_{3}(k)$, initially draw them disjointly as previously described, and then modify and interleave the three drawings so as to contain no [Uncaptioned image]-cell. See Figure˜12 for an illustration. Index operations for the $G^{\prime}_{i}(k)$ will be performed modulo 3. We start by placing the three copies in three disjoint disks, which we call the centers of the respective $G^{\prime}_{i}(k)$. We shall move the vertices of $G^{\prime}_{i}(k)$ in such a way that each vertex $a$ of $G^{\prime}_{i}(k)$ lies close to a point $b$ in the set $B$ of $G^{\prime}_{i+1}(k)$. We locally perturb the $k$ edges of $G^{\prime}_{i+1}(k)$ that all cross in $b$ so that each of the $k$ edges of $G^{\prime}_{i}(k)$ emanating from $a$ has its first crossing with a different edge of $G^{\prime}_{i+1}(k)$. Loosely speaking, the edges of $G^{\prime}_{i+1}(k)$ at point $b$ “support” the edges of $G^{\prime}_{i}(k)$ starting at point $a$.
@@
More precisely, we consider in the drawing of $G^{\prime}(k)$ for each $b\in B$ a straight and very narrow corridor starting at $b$ and ending on $C_{\rm out}$ between $a_{0}$ and $a_{3k-2}$. These corridors are routed so that they cross no existing edge twice. Then, we perturb the $k$ edges close to $b$ to obtain a simple drawing in which $b$ sees a segment of each of these $k$ edges. Now, place the simple drawings of $G^{\prime}_{1}(k),G^{\prime}_{2}(k),G^{\prime}_{3}(k)$ with corridors disjointly next to each other. Move the vertices of $G^{\prime}_{1}(k)$ (without introducing new crossings between edges of $G^{\prime}_{1}(k)$) through the prepared corridors of $G^{\prime}_{2}(k)$ (each vertex through one corridor) into the center of $G^{\prime}_{2}(k)$. Place each vertex $a$ of $G^{\prime}_{1}(k)$ at the end of the corridor of $G^{\prime}_{2}(k)$ (at the point $b\in B$) so that each edge (of $G^{\prime}_{1}(k)$) at $a$ crosses through a different edge (of $G^{\prime}_{2}(k)$) near $b$. This ensures that $a$ is not part of any [Uncaptioned image]-cell. By similarly threading the vertices of $G^{\prime}_{2}(k)$ through the corridors of $G^{\prime}_{3}(k)$ and the vertices of $G^{\prime}_{3}(k)$ through the corridors of $G^{\prime}_{1}(k)$, [Uncaptioned image]-freeness is ensured. Note that every edge of $G^{\prime}_{i}(k)$ now passes exactly two vertices of $G^{\prime}_{i-1}(k)$ that are now placed within the center of $G^{\prime}_{i}(k)$. This finishes the construction of the drawing of $G(k)$.
@@
We are left with arguing non-homotopicity. First note that, while each component is itself drawn simply and no corridor crosses any edge twice, the resulting drawing is not simple. The reason for this is that every edge passes through two corridors (near its two incident vertices), which may cross a common set of edges (see again Figure˜12). More precisely, all violations to simplicity occur because an edge $f$ of $G_{i}^{\prime}(k)$ that starts and ends in the center of $G_{i+1}(k)$ is crossed twice by an edge $f^{\prime}$ of $G_{i+1}(k)$ within the center of $G_{i+1}(k)$. To argue that the drawing is non-homotopic, we show that any such lens $\ell$ bounded by $f$ and $f^{\prime}$ is non-empty. Note that, by construction, $f$ passes near two vertices $u$ and $v$ of $G_{i-1}^{\prime}(k)$ within the center of $G_{i}(k)$. We will show that $u$ and $v$ lie inside the lens $\ell$. For contradiction’s sake, assume that one of the two vertices, say $u$, was not contained in the interior of $\ell$. Again, by construction, $f$ “supports“ an edge $g$ incident to $u$, that is, $g$ has its first crossing with $f$, where by assumption it enters $\ell$. After this crossing, the edge $g$ has to follow a corridor to enter the center of $G_{i-1}^{\prime}(k)$. However, outside the center of $G_{i}^{\prime}(k)$, the lens $\ell$ does not contain any part of a corridor connecting the center of $G_{i}^{\prime}(k)$ to the center of $G_{i-1}^{\prime}(k)$. Hence, $g$ needs to cross $f$ again within the center of $G_{i-1}^{\prime}(k)$. But this contradicts the fact that every corridor is drawn simply with respect to the three initial copies of $G^{\prime}(k)$. This contradiction shows that $\ell$ contains $u$ (and also $v$) in its interior. The same argument holds for every lens in the drawing, and thus, our drawing of $G(k)$ is non-homotopic.

### v3

@@ 5.2 Simple Non-Quasiplanar [Uncaptioned image]-Free Drawings
There are infinitely many graphs that admit a simple [Uncaptioned image]-free drawing but no non-homotopic quasiplanar drawing.
Proof 5.5.
Start with an arbitrary graph $G$ that does not admit a non-homotopic quasiplanar drawing (any graph on $n$ vertices and >8n20>8n-20 edges suffices). Let $\Gamma$ be any simple drawing of $G$. If $\Gamma$ contains no [Uncaptioned image]-cells, we are done. Otherwise, let $c$ be a [Uncaptioned image]-cell and $e=uv$ one of the edges on its boundary. Add a new vertex $w$ in $c$ to $\Gamma$ and draw an edge $uw$ that starts at $u$, follows $e$ closely (such that $uw$ crosses the exact same edges as $e$ does between $u$ and the part where $e$ bounds the [Uncaptioned image]-cell containing $w$, in the same order) until entering $c$ and ending in $w$; see Figure˜7. In the new drawing $\Gamma^{+}$, the corridor bounded by $e$ and $uw$ only contains a [Uncaptioned image]-cell and some [Uncaptioned image]-cells. Moreover, $c$ becomes a (degenerate) [Uncaptioned image]-cell in $\Gamma^{+}$. All other cells in $\Gamma^{+}$ incident to $uw$ have a corresponding cell of the same type in $\Gamma$ with the only difference that the edge segments of $e$ are replaced by edge segments of $uw$. Therefore, adding $w$ and $uw$ reduces the number of [Uncaptioned image]-cells. Further, the new drawing $\Gamma^{+}$ retains simplicity, as $uw$ crosses a subset of the edges crossed by $e$.
Repeating this process, a simple [Uncaptioned image]-free drawing $\Gamma^{\prime}$ of a graph $G^{\prime}$ with $G\subseteq G^{\prime}$ is obtained. Since $G$ does not admit a non-homotopic quasiplanar drawing, neither does $G^{\prime}$, and so, $G^{\prime}$ is as desired.
Figure 7: Illustration of Proposition˜5.4. By adding a vertex and an edge, the number of Refer to caption-cells in a drawing can be reduced by one.
@@ 6.1 Simple [Uncaptioned image]-Free Drawings
We begin with the construction of simple [Uncaptioned image]-free drawings on the torus with a quadratic number of edges.
@@
Now, we add for each vertex $v$ in column $j$ and each of the $k$ slopes $s\in S_{j}$ an edge with slope $s$ from $v$ to a vertex in column $j+3$. That is, for each vertex $v_{i,j}$, and for each slope $s\in S_{j}$, we add an edge with slope $s$ between $v_{i,j}$ and $v_{i+3s,j+3}$ that passes through $v_{i+s,j+1}$ and $v_{i+2s,j+2}$. See Figure˜8. Note that this is not yet a valid drawing by our definition, as edges pass through vertices and there are points in which more than two edges cross.
To fix this, consider any vertex $v_{i,j}$. There are $k$ edges incident to $v_{i,j}$ that end in column $j+3$, and $k$ edges incident to $v_{i,j}$ that start in column j3j-3. Since $S_{j}=S_{j-3}=S_{j+3}$, both these sets of $k$ edges have the same $k$ slopes. Additionally, there is a set $A$ of $2k$ edges that pass through $v_{i,j}$ but are not incident to $v_{i,j}$. The $2k$ edges in $A$ have the $2k$ slopes in $\{0,\ldots,3k-1\}-S_{j}$. The situation in a small disk around $v_{i,j}$ is depicted in Figure˜9 (left). We perturb the $2k$ edges of $A$ slightly within the disk around $v_{i,j}$ such that locally we obtain a valid drawing with the additional property that each of the $2k$ edges emanating from $v_{i,j}$ has its first crossing with a different of the $2k$ edges in $A$ (see Figure˜9 (right)). This ensures that none of the cells incident to $v_{i,j}$ is a [Uncaptioned image]-cell. Repeating this procedure at every vertex $v_{i,j}$, the resulting drawing is [Uncaptioned image]-free. At each remaining point (in the space between the columns) in which more than two edges cross, we similarly locally perturb the edges to eventually obtain a valid drawing $\Gamma(k)$.
Figure 9: Illustration for Construction˜9. In a drawing, when a set $A$ of $2k$ edges intersect at a vertex of degree $2k$, the edges can be locally redrawn without introducing new edge intersections such that the vertex is not incident to any Refer to caption-cells.
Finally, we argue that $\Gamma(k)$ is a simple drawing. Note that any edge in $\Gamma(k)$ spans (including its endpoints) at most $9k$ rows and exactly $4$ columns. Hence the size of the grid ensures that two edges do not intersect multiple times by wrapping around the torus. Further, before the perturbations, all edges locally behave like straight-line segments, and hence any two (unperturbed) edges intersect at most once. Because the perturbations do not introduce any new crossings, $\Gamma(k)$ is simple.
Based on the toroidal construction, we obtain simple [Uncaptioned image]-free drawings in the plane with $\Omega(n^{3/2})$ edges. While this leaves a significant gap to the trivial upper bound of $\binom{n}{2}$, we believe that there are [Uncaptioned image]-free drawings with $\Omega(n^{2})$ edges, see ˜8.2.
@@
Proof 6.2.
Let $k$ be some positive integer. Start with the toroidal drawing $\Gamma(k)$ described in Construction˜9 but with an underlying $20k\times 9k$-grid instead of a $20k\times 9$-grid. See again Figure˜8 for a visualization. Now, since we cannot place $\Gamma(k)$ in the plane as is, we cut the torus between two neighboring columns (say between column 9k19k-1 and 0). The cut is positioned so that it does not pass through any edge crossings. The resulting “drawing” can be embedded in the plane but contains many halved edges that end on the boundaries of the cut we performed. We take care of these half-edges by placing a new vertex each at the point in the plane where the half-edge ends, and, in doing so, add degree 1-vertices.
@@ 6.2 Non-Homotopic [Uncaptioned image]-Free Drawings
For each integer $k\geq 2$, we construct a simple $k$-regular graph $G(k)$ on $n=3\cdot(3k-1)$ vertices together with a non-homotopic [Uncaptioned image]-free drawing; see Figures˜10 and 11 for an illustration. The graph $G(k)$ consists of three copies of a graph $G^{\prime}(k)$ on 3k13k-1 vertices. We describe $G^{\prime}(k)$ in tandem with a temporary ?drawing? that does not yet fulfill our definition of a drawing. We start with a set $A$ of 3k13k-1 equidistant points on a circle $C_{\rm out}$. The points in $A$ will be the vertices of $G^{\prime}(k)$, which we denote $a_{0},\ldots,a_{3k-2}$ in cyclic ordering around $C_{\rm out}$. For each $i=0,\ldots,3k-2$ draw a straight-line edge $e_{i}$ between $a_{i}$ and $a_{i+k}$ (throughout this construction all indices are taken modulo 3k13k-1). Let $b_{i}$ be the crossing point of $e_{i}$ and $e_{i+1}$, and $B$ the set of all such $b_{i}$.
@@
Figure 10: Illustration of the $k$-regular graph $G^{\prime}(k)$ on 3k13k-1 vertices for $k=2$ and $k=5$. At each of the 3k13k-1 points in $B$ (blue disks), $k$ edges are pairwise crossing. The zoomed-in subfigure depicts the situation at a point $b\in B$ after interleaving the three copies of $G^{\prime}(k)$ to obtain $G(k)$ (see also Figure˜11). In particular, the red elements (including the corridors) belong to a different copy of $G^{\prime}(k)$. For visual clarity, only some corridors are shown for $k=5$.
@@
Then $b_{0},\ldots,b_{3k-2}$ are equidistant points (blue disks in Figure˜10) on a smaller circle $C_{\rm in}$ concentric to $C_{\rm out}$. Note that $e_{i}$ crosses $e_{i+1}$ at $b_{i}$ and $e_{i-1}$ at $b_{i-1}$. Next, we add further edges such that each $a_{i}$ has exactly $k$ incident edges and at each $b_{i}$, exactly $k$ edges pairwise cross (thus making the result not a drawing in our sense). To this end, we insert edges from $a_{i}$ to $a_{i+k+1},\ldots,a_{i+\lfloor 3k/2\rfloor}$ where the edge $a_{i}a_{i+k+j}$ with $j=1,\ldots,\lfloor k/2\rfloor$ goes through $b_{i-j-1}$ and $b_{i+2j}$; see again Figure˜10 for an illustration. Let us emphasize (as we will need it later) that every edge passes through exactly two points in $B$. This completes the construction of $G^{\prime}(k)$. Observe that the described drawing of $G^{\prime}(k)$ is simple (in fact, almost straight-line) and every vertex of $G^{\prime}(k)$ lies on the same cell.
@@
Figure 11: Connecting the drawings of $G^{\prime}_{1}(k)$, $G^{\prime}_{2}(k)$, and $G^{\prime}_{3}(k)$ to a Refer to caption-free drawing of $G(k)$. The center of each $G^{\prime}_{i}(k)$ is depicted as a yellow disk. The two edges $f$ and $f^{\prime}$ cross twice (marked orange) but the lens they create is non-empty, since it contains the vertices $u$ and $v$.
@@
We now take three copies of $G^{\prime}(k)$, called $G^{\prime}_{1}(k),G^{\prime}_{2}(k),G^{\prime}_{3}(k)$, initially draw them disjointly as previously described, and then modify and interleave the three ?drawings? so as to obtain a valid drawing that contains no [Uncaptioned image]-cell. See Figure˜11 for an illustration. Index operations for the $G^{\prime}_{i}(k)$ will be performed modulo 3. We start by placing the three copies in three disjoint disks, which we call the centers of the respective $G^{\prime}_{i}(k)$. We shall move the vertices of $G^{\prime}_{i}(k)$ in such a way that each vertex $a$ of $G^{\prime}_{i}(k)$ lies close to a point $b$ in the set $B$ of $G^{\prime}_{i+1}(k)$. We locally perturb the $k$ edges of $G^{\prime}_{i+1}(k)$ that all cross in $b$ so that each of the $k$ edges of $G^{\prime}_{i}(k)$ emanating from $a$ has its first crossing with a different edge of $G^{\prime}_{i+1}(k)$, see the zoomed-in subfigure in Figure˜10. Loosely speaking, the edges of $G^{\prime}_{i+1}(k)$ that were crossing in a point $b$ before being perturbed “support” the edges of $G^{\prime}_{i}(k)$ starting at a vertex $a$ of $G^{\prime}_{i}(k)$.
@@
More precisely, we consider in the temporary drawing of $G^{\prime}(k)$ for each $b\in B$ a straight and very narrow corridor starting at point $b$ and ending on $C_{\rm out}$ between $a_{0}$ and $a_{3k-2}$. These corridors are routed so that they intersect no existing edge twice (including those currently passing through $b$). We perturb the $k$ edges close to $b$ to obtain a simple drawing in which $b$ sees a segment of each of these $k$ edges. This is similar to the redrawing procedure depicted in Figure˜9 and ensures that the final drawing is valid. Now, place the simple drawings of $G^{\prime}_{1}(k),G^{\prime}_{2}(k),G^{\prime}_{3}(k)$ with corridors disjointly next to each other. Move the vertices of $G^{\prime}_{1}(k)$ (without introducing new crossings between edges of $G^{\prime}_{1}(k)$) through the prepared corridors of $G^{\prime}_{2}(k)$ (each vertex through one corridor) into the center of $G^{\prime}_{2}(k)$. Place each vertex $a$ of $G^{\prime}_{1}(k)$ at the end of the corridor of $G^{\prime}_{2}(k)$ (at the point $b\in B$) so that each edge (of $G^{\prime}_{1}(k)$) at $a$ crosses through a different edge (of $G^{\prime}_{2}(k)$) near $b$. This ensures that $a$ is not part of any [Uncaptioned image]-cell. By similarly threading the vertices of $G^{\prime}_{2}(k)$ through the corridors of $G^{\prime}_{3}(k)$ and the vertices of $G^{\prime}_{3}(k)$ through the corridors of $G^{\prime}_{1}(k)$, [Uncaptioned image]-freeness is ensured. Note that every edge of $G^{\prime}_{i}(k)$ now passes exactly two vertices of $G^{\prime}_{i-1}(k)$ that are now placed within the center of $G^{\prime}_{i}(k)$. This finishes the construction of the drawing of $G(k)$.
@@
We are left with arguing non-homotopicity. First note that, while each component is itself drawn simply and no corridor crosses any edge twice, the resulting drawing is not simple. The reason for this is that every edge passes through two corridors (near its two incident vertices), which may cross a common set of edges (see again Figure˜11). More precisely, all violations to simplicity occur because an edge $f$ of $G_{i}^{\prime}(k)$ that starts and ends in the center of $G^{\prime}_{i+1}(k)$ is crossed twice by an edge $f^{\prime}$ of $G^{\prime}_{i+1}(k)$ within the center of $G^{\prime}_{i+1}(k)$. To argue that the drawing is non-homotopic, we show that any such lens $\ell$ bounded by $f$ and $f^{\prime}$ is non-empty. Note that, by construction, $f$ passes near two vertices $u$ and $v$ of $G_{i-1}^{\prime}(k)$ within the center of $G^{\prime}_{i}(k)$. We will show that $u$ and $v$ lie inside the lens $\ell$. For contradiction’s sake, assume that one of the two vertices, say $u$, was not contained in the interior of $\ell$. Again, by construction, $f$ “supports“ an edge $g$ incident to $u$. That is, $g$ has its first crossing with $f$, where by assumption it enters $\ell$. Because of the way the copies are interleaved, $g$ has to follow a corridor to enter the center of $G_{i-1}^{\prime}(k)$ after crossing $f$. However, none of the corridors that connect the center of $G_{i}^{\prime}(k)$ to the center of $G_{i-1}^{\prime}(k)$ are contained in the interior of $\ell$, except for their initial segments inside the center of $G_{i}^{\prime}(k)$. Hence, in order to follow a corridor into the center of $G_{i-1}(k)$, $g$ needs to cross $f$ again within the center of $G_{i}^{\prime}(k)$. But then the corridor through which $g$ travels crosses $f$ twice, a contradiction to the fact that the corridors are drawn simply with respect to the initial drawings of $G^{\prime}(k)$. This contradiction shows that $\ell$ contains $u$ (and also $v$) in its interior. The same argument holds for every lens in the drawing, and thus, our drawing of $G(k)$ is non-homotopic.